dave wants to purchase 21 pounds of party mix for a total of $58. to obtain the mixture, he will mix nuts that cost $4 per pound woth pretzels that cost $2 per pound. How many pounds of each type and mix should he use?
Answer by addingup(3677)

n+p = 21, so n = 21-p
4n+2p = 58
4(21-p)+2p = 58
84-4p+2p = 58
-2p = -26 Divide both sides by -2 and remember that -/- = +
p = 13 this is the pretzels, and the mix nuts:
n= 21-13 = 8